OpenStudy (anonymous):

QUESTION 5.... Out of 10 2x+5y=15 5x+10y=35 A. (10, -1) B. (5, 1) C. (5, -1) D. (10, 1)

OpenStudy (mrnood):

Please do not give direct answers - if there are 10 questions then guidance should help the OP to solve the rest

OpenStudy (anonymous):

lets use the good old elimination rule. again. multiply first equation by -2 -4x -10y = -30 5x + 10y = 35 sum them up and get x = 5 2 * 5 + 5y = 15 10 + 5y = 15 5y = 5 y = 1 Check: 25 + 10 = 35 35 = 35 Variant B(5 1)

OpenStudy (mrnood):

= direct answer

OpenStudy (anonymous):

not at all. at first i gave full explanation. wxplaining the elimination method already for around 4 questions.

OpenStudy (mrnood):

So why not let OP use your help and work through the process for herself for the next 5?

OpenStudy (mrnood):

All she will take from your answer is "B"

OpenStudy (anonymous):

At first she will read the explanation. there are things that you alwasy fail at? i have such things like probability theory and discrete math and i fail them all the time. People here help me at them. however if you want you can throw rock at me.